Show me how to do this problem sum of two cubes and show work
x^3+27

Answer by mananth(16949)

The sum of cubes identity is
a^3+b^3= (a+b)(a^2-ab+b^2)
x^3 +27
x^3+3^3
here a=x and b=3
(x+3)(x^2-3x+9)